Use the Substitution method to solve the system of equations. Give your answer as an ordered pair (x,y).


x + y = -5
x - y = 3

Answer by funmath(2933)

Use the Substitution method to solve the system of equations. Give your answer as an ordered pair (x,y).


E1: x + y = -5
E2: x - y = 3
Solve E2 for x.
x-y+y=3+y
x=3+y
Substitute 3+y into E1 for x and solve for y
(3+y)+y=-5
3+2y=-5
3-3+2y=-5-3
2y=-8
2y/2=-8/2
y=-4
Substitute -4 in for y in E2 and solve for x
x-(-4)=3
x+4=3
x+4-4=3-4
x=-1
The solution (x,y)=(-1,-4)
You can check your answer by substituting the solution into both E1 and E2 and see if both sides equal.
-1-4=-5 --> -5=-5 E1 checks
-1-(-4)=3 --> 3=3 E2 check too.
Happy Calculating!!!